# ProDAG: Projection-Induced Variational Inference for Directed Acyclic Graphs

### v4

@@ Abstract
Directed acyclic graph (DAG) learning is a rapidly expanding field of research. Though the field has witnessed remarkable advances over the past few years, it remains statistically and computationally challenging to learn a single (point estimate) DAG from data, let alone provide uncertainty quantification. Our paper addresses the difficult task of quantifying graph uncertainty by developing a Bayesian variational inference framework based on novel distributions that have support directly on the space of DAGs. The distributions, which we use to form our prior and variational posterior, are induced by a projection operation, whereby an arbitrary continuous distribution is projected onto the space of sparse weighted acyclic adjacency matrices (matrix representations of DAGs) with probability mass on exact zeros. Though the projection constitutes a combinatorial optimization problem, it is solvable at scale via recently developed techniques that reformulate acyclicity as a continuous constraint. We empirically demonstrate that our proposed method, ProDAG, can perform higher quality Bayesian inference than possible with existing state-of-the-art alternatives.
@@ 1 Introduction
A directed acyclic graph (DAG) has directed edges between nodes and no directed cycles, allowing it to represent complex conditional independencies compactly. DAGs play a central role in causal inference (Spirtes et al., 2001; Pearl, 2009) and consequently have found applications spanning a broad range of domains such as psychology (Foster, 2010), economics (Imbens, 2020), and epidemiology (Tennant et al., 2021). The literature on DAGs, dating back at least to the 1980s (Lauritzen & Spiegelhalter, 1988; Pearl, 1988), has recently witnessed a resurgence of interest due to a computational breakthrough by Zheng et al. (2018) that formulated a continuous characterization for the discrete notion of acyclicity. This breakthrough opened the door for scalable first-order algorithms that can learn DAG structures from datasets containing many more variables than previously assumed feasible. We refer the reader to Vowels et al. (2022) and Kitson et al. (2023) for surveys of this work.
@@
Our paper introduces ProDAG, a Bayesian method for learning DAGs that departs from the approaches above. Leveraging the benefits of variational inference, ProDAG models the prior and variational posterior using new distributions that directly have support on the space of DAGs. These DAG distributions are induced by mapping a continuous distribution $\mathbb{P}$ onto the space of DAGs using a minimal distance projection. Consider a DAG represented by the weighted adjacency matrix $W\in\mathbb{R}^{p\times p}$ with entries $\mathsf{w}_{jk}$ that are nonzero if and only if there exists a directed edge from node $j$ to node $k$. We induce a distribution over such matrices $W$, and hence over DAGs, using the data generating process
@@
where $\lambda\geq 0$ is a parameter governing the graph’s sparsity (number of active/nonzero edges) with prior $\pi$ and
@@
is the projection of the continuous matrix $\tilde{W}\in\mathbb{R}^{p\times p}$ onto the set of $\ell_{1}$ constrained acyclic matrices. [^2] The norm $\|W\|_{\ell_{1}}:=\|\operatorname{vec}(W)\|_{1}$. We prove this projection is a measurable mapping with a unique solution, yielding valid probability distributions. One of the key qualities of these distributions is that they place probability mass on exact zeros, a necessity for sampling DAGs.
Inspired by recently developed continuous DAG characterizations (e.g., Zheng et al., 2018), we reformulate the combinatorial optimization problem (2) as an equivalent continuous optimization problem that can be solved in parallel on a GPU. We use this formulation in a variational framework by forming the prior and variational posterior using the new distributions, enabling scalable Bayesian DAG learning. Our framework accommodates DAGs represented as linear and nonlinear structural equation models (SEMs), making our approach suitable for problems and data of varying complexity. Experiments on synthetic and real data demonstrate that our approach often delivers more accurate inference than existing methods for DAG learning. Our toolkit implementing ProDAG is released open-source.
@@ 2 Related Work
As mentioned, existing approaches for Bayesian inference on DAGs broadly fall into two groups: (1) regularization-based methods that include a continuous acyclicity penalty via a Gibbs-type prior in the variational objective (Annadani et al., 2021; Lorch et al., 2021; Geffner et al., 2022), and (2) state-augmentation methods that, e.g., place a prior over the graph’s topological ordering (Cundy et al., 2021; Charpentier et al., 2022; Bonilla et al., 2024; Toth et al., 2024). The node-potential augmentation of Annadani et al. (2023a) is related to the latter category. See also Wang et al. (2022); Deleu et al. (2023) for approaches based on probabilistic circuits and generative flow networks. Our approach differs by designing a prior and approximate posterior directly over DAGs without modeling the topological ordering and without using a penalty that does not place mass on zeros.
More broadly, our paper complements the literature on computational methods for point estimate DAG learning. The seminal paper of Zheng et al. (2018), which introduced the NOTEARS method, cleverly exploited the correspondence between the number of cycles in a matrix and the trace of that matrix’s power. A similar connection is now known to hold for other continuous functions and is exploited by methods such as DAGMA (Bello et al., 2022), which we utilize in this work, as well as those of Ng et al. (2020), Yu et al. (2021), and Gillot & Parviainen (2022). These approaches that use continuous characterizations of acyclicity are appealing since they scale to large graphs and can be augmented with discrete combinatorics (Manzour et al., 2021; Deng et al., 2023) if required. Though not pursued here, our DAG projection is also amenable to discrete techniques.
Finally, though distributions over DAGs via projection is a new idea, others have used projections to construct distributions over non-graphical structures. Xu & Duan (2023) designed distributions over sparse regression coefficients by projecting a vector drawn from an arbitrary continuous distribution onto the $\ell_{1}$ ball. They illustrated several computational advantages of their distributions as priors in a Hamiltonian Monte Carlo (HMC) framework over classical spike-and-slab priors and prove that the resulting posterior enjoys the minimax concentration rate. Xu et al. (2023) extended this work to general proximal mappings with HMC, which includes the convex $\ell_{1}$ projection as a special case. While our approach projects matrices onto the $\ell_{1}$ ball, we also impose acyclicity, giving rise to a nonconvex combinatorial optimization problem that is not a proximal mapping.
@@ 3.1 Description
To recap, we induce a distribution over weighted acyclic adjacency matrices $W$ as $W=\operatorname{pro}_{\lambda}(\tilde{W})$, where $\operatorname{pro}_{\lambda}$ is defined in (2), $\tilde{W}\sim\mathbb{P}$ is a continuous matrix, and $\lambda\sim\pi$ bounds the $\ell_{1}$ norm of $W$. Since acyclic graphs have no self-loops (i.e., $W$ always has a zero diagonal), we can take $\mathbb{P}$ to be any continuous distribution over the off-diagonal elements of $\tilde{W}$ (e.g., a multivariate Gaussian), with the diagonal elements fixed as zero. The matrix $\tilde{W}$ can thus be interpreted as the weighted adjacency matrix of a simple directed graph, with the projection removing all cycles. [^3] Simple directed graphs have no self-loops and, at most, one edge in each direction between any pair of nodes.
Besides removing cycles, the projection maps $\tilde{W}$ onto an $\ell_{1}$ ball of size $\lambda$. This part of the projection is necessary for sampling parsimonious graphs. Without this constraint, the projection would return the densest possible acyclic graph with $(p^{2}-p)/2$ edges. [^4] An exception is the degenerate case where some off-diagonal elements of $\tilde{W}$ are already zero. The advantage of constraining $W$ to an $\ell_{1}$ ball, compared with modeling $\tilde{W}$ as a Laplace distribution, is that the $\ell_{1}$ ball gives probability mass on exact zeros beyond that already given by the acyclicity constraint. It is known that Laplace distributions do not concentrate exactly on zero except at their mode (Park & Casella, 2008).
The resulting distribution $p(W)$ can be understood as a mix of continuous components and exact zeros. To elaborate, first observe that the nonzeros in $W$ may be indexed by a sparsity pattern $S\in\{0,1\}^{p\times p}$ (i.e., an unweighted acyclic adjacency matrix). The nonzeros are then given by $W_{S}$, where $W_{S}$ is the restriction of $W$ to $S$. Because the $\ell_{1}$ constraint shrinks the nonzeros towards zero, the continuous $W_{S}$ is a transformation $f_{\lambda}$ of $\tilde{W}_{S}$, i.e., $W_{S}=f_{\lambda}(\tilde{W}_{S})$, where $f_{\lambda}$ is implicit in the projection. This mixture interpretation suggests an interesting representation of $p(W)$:
$\begin{split}p(W)&\equiv\sum_{i=1}^{m}p(W_{S_{i}}\mid S_{i})p(S_{i})\\ &=\sum_{i=1}^{m}p(f_{\lambda}(\tilde{W}_{S_{i}})\mid S_{i})p(S_{i}),\end{split}$
where $S_{1},\dots,S_{m}$ is the finite collection of all acyclic adjacency matrices. To reiterate, the distributions $p(f_{\lambda}(\tilde{W}_{S})\mid S)$ and $p(S)$ on the right-hand side are implicitly defined by the projection. Though this representation is conceptual, it highlights that $p(W)$ forms a distribution over acyclic adjacency matrices $S$ without explicitly having to model the discrete, high-dimensional probability mass function $p(S)$. Moreover, as we later show, the projection can be reformulated as a continuous optimization problem, meaning one can sample from $p(W)$ without combinatorial complexity.
@@ 3.2 Properties
One might ask if $p(W)$ is a valid probability distribution. Theorem 3.1 addresses this question in the affirmative by showing that the projection is almost surely unique and measurable. The proof of the theorem is in Appendix A.
Theorem 3.1.
@@
Theorem 3.1 states that, for any continuous $\tilde{W}\sim\mathbb{P}$, the function $\operatorname{pro}_{\lambda}(\tilde{W})$ is uniquely defined and measurable. This result implies that a projected distribution is a valid distribution since a measurable function $g$ applied to a random variable $z$ produces a well-defined random variable $g(z)$.
3.3 Graph Sparsity
To attain a distribution over graphs that contain a similar number of edges, the distribution $\pi$ for the parameter $\lambda$ can be configured as a Dirac delta distribution concentrated on a point $\alpha\geq 0$ that yields the desired level of sparsity:
$\pi_{\alpha}(\lambda)=\delta(\lambda-\alpha).$
Alternatively, for graphs of varying complexity, $\pi$ can be specified as a distribution over $\lambda\geq 0$. For instance, $\pi$ can be taken as an exponential distribution with mean $\alpha\geq 0$:
$\pi_{\alpha}(\lambda)=\frac{\exp(-\lambda/\alpha)}{\alpha},\quad\lambda\geq 0.$
This specification allows for graphs with varying numbers of edges but primarily favors sparser (parsimonious) graphs over denser graphs for smaller values of $\alpha$. The mean $\alpha$ of $\lambda$ can be determined using expert domain knowledge or empirically calibrated on an available validation set.
@@ 4.1 Continuous Reformulation
To facilitate scalability, we reformulate the combinatorial projection (2) as an equivalent continuous optimization problem. This recharacterization follows recent work (Zheng et al., 2018; Bello et al., 2022) that established an an equivalence relationship for certain functions $h(W)$:
$h(W)=0\Longleftrightarrow W\in\mathrm{DAG}.$
Bello et al. (2022) showed that for all matrices $W\in\mathbb{W}:=\{W\in\mathbb{R}^{p\times p}:\rho(W\circ W)<1\}$, where $\rho$ is the spectral radius and $\circ$ is the Hadamard product, the function $h(W):=-\log\,\det(I-W\circ W)$ satisfies this property. This function, which constitutes the main innovation of DAGMA, attains the value zero if and only if there are no cycles in $W$. Thus, its level set at zero exactly corresponds to the entire set of DAGs. Though it is a nonconvex function, $h(W)$ is useful because it is both continuous and differentiable in $W$.
We now rewrite projection (2) as the equivalent projection
@@
As we discuss next, it is possible to (a) solve this projection using only first-order methods and (b) analytically evaluate the gradients of this projection. In turn, these attributes pave the way for scalable, gradient-based posterior learning.
@@ 4.2 Algorithms
To solve projection (3), we employ a two-step procedure that produces a solution on the intersection of sets. Step 1 of this procedure projects the continuous matrix $\tilde{W}$ onto the set of acyclic matrices. We refer to the result of Step 1 as $\hat{W}$. Step 2 projects $\hat{W}$ onto the $\ell_{1}$ ball, yielding the final output matrix $W$ that remains acyclic but is further sparse.
For the acyclicity projection in the first step, we use a path-following algorithm similar to Bello et al. (2022). This algorithm reformulates the constrained problem as a sequence of unconstrained problems involving a penalized objective:

### v5

# ProDAG: Projected Variational Inference for Directed Acyclic Graphs

@@ Abstract
Directed acyclic graph (DAG) learning is a central task in structure discovery and causal inference. Although the field has witnessed remarkable advances over the past few years, it remains statistically and computationally challenging to learn a single (point estimate) DAG from data, let alone provide uncertainty quantification. We address the difficult task of quantifying graph uncertainty by developing a Bayesian variational inference framework based on novel, provably valid distributions that have support directly on the space of sparse DAGs. These distributions, which we use to define our prior and variational posterior, are induced by a projection operation that maps an arbitrary continuous distribution onto the space of sparse weighted acyclic adjacency matrices. While this projection is combinatorial, it can be solved efficiently using recent continuous reformulations of acyclicity constraints. We empirically demonstrate that our method, ProDAG, can outperform state-of-the-art alternatives in both accuracy and uncertainty quantification.
@@ 1 Introduction
A directed acyclic graph (DAG) has directed edges between nodes and no directed cycles, allowing it to represent complex conditional independencies compactly. DAGs play a central role in causal inference (Spirtes et al., 2001; Pearl, 2009) and consequently have found applications spanning a broad range of domains such as psychology (Foster, 2010), economics (Imbens, 2020), and epidemiology (Tennant et al., 2021). The literature on DAGs, dating back at least to the 1980s (Lauritzen and Spiegelhalter, 1988; Pearl, 1988), has recently witnessed a resurgence of interest due to a computational breakthrough by Zheng et al. (2018) that formulated a continuous characterization for the discrete notion of acyclicity. This breakthrough opened the door for scalable first-order algorithms that can learn DAG structures from datasets containing many more variables than previously assumed feasible. We refer the reader to Vowels et al. (2022) and Kitson et al. (2023) for surveys of this work.
@@
We propose ProDAG, a Bayesian method for learning DAGs that departs from the approaches above. Leveraging the power of variational inference, ProDAG models the prior and variational posterior using new distributions that natively have support on the space of DAGs. These DAG distributions are induced by mapping a continuous distribution $\mathbb{P}$ onto the space of DAGs via a minimal distance projection. Specifically, consider a DAG represented by the weighted adjacency matrix $W\in\mathbb{R}^{p\times p}$ with entry $\mathsf{w}_{jk}$ nonzero if and only if a directed edge exists from node $j$ to node $k$. We define a distribution over such matrices $W$, and hence over DAGs, via the data generating process
@@
where $\lambda\geq 0$ is a parameter governing the graph’s sparsity (number of edges) with prior $\pi$ and
@@
is the projection of the continuous matrix $\tilde{W}\in\mathbb{R}^{p\times p}$ onto the set of $\ell_{1}$ constrained acyclic matrices. Figure 1 visually illustrates the process.
Figure 1: Illustration of ProDAG’s projected distributions. Samples (blue dots) from an unconstrained continuous space (blue ellipse) are projected onto the nearest acyclic matrix within an $\ell_{1}$-constrained region (orange diamonds). Projected samples (orange dots) satisfy acyclicity and sparsity constraints. Theoretically, we show that for any continuous $\tilde{W}\sim\mathbb{P}$ the function $\operatorname{pro}_{\lambda}(\tilde{W})$ is uniquely defined and measurable. This implies that the projected distribution is a valid distribution over DAGs.
We prove this projection is a measurable mapping with a unique solution, yielding valid probability distributions. A key quality of these distributions is that they place probability mass on exact zeros, necessary for sampling valid, sparsely connected DAGs.
Inspired by recently developed continuous DAG characterizations (e.g., Zheng et al., 2018), we reformulate the combinatorial optimization problem (2) as an equivalent continuous optimization problem that can be solved in parallel on a GPU. We use this formulation in a variational framework by forming the prior and variational posterior using the new distributions, enabling scalable Bayesian DAG learning. Our framework accommodates DAGs represented as linear and nonlinear structural equation models (SEMs), making our approach suitable for problems and data of varying complexity. Experiments on synthetic and real data demonstrate that our approach often delivers more accurate inference than existing methods for DAG learning. Our toolkit for ProDAG is available on GitHub.
In summary, our core contributions are:
  1. 1.
    A new Bayesian model that directly samples from provably valid DAG distributions, ensuring exact acyclicity and eliminating the need for discrete approximations or relaxations;
  2. 2.
    A scalable variational inference framework that leverages GPU-accelerated projections, enabling efficient and accurate uncertainty quantification for linear and nonlinear DAGs;
  3. 3.
    An extensive suite of empirical evaluations that demonstrate state-of-the-art results across a wide variety of synthetic datasets, and validated on a real biological dataset; and
  4. 4.
    A user-friendly, well-documented, and open-source Julia implementation, made publicly available to promote adoption and reproducibility among researchers and practitioners.
@@ 2 Related work
As mentioned, existing approaches for Bayesian inference on DAGs broadly fall into two groups: (1) regularization-based methods that include a continuous acyclicity penalty via a Gibbs-type prior in the variational objective (Annadani et al., 2021; Lorch et al., 2021; Geffner et al., 2022), and (2) state-augmentation methods that, e.g., place a prior over the graph’s topological ordering (Cundy et al., 2021; Charpentier et al., 2022; Bonilla et al., 2024; Toth et al., 2024). The node-potential augmentation of Annadani et al. (2023a) is related to the latter category. See also Wang et al. (2022); Deleu et al. (2023) for approaches based on probabilistic circuits and generative flow networks. Our approach differs by designing a prior and approximate posterior directly over DAGs without modeling the (discrete) topological ordering and without using a penalty that does not place mass on zeros.
More broadly, our paper complements the literature on computational methods for point estimate DAG learning. The seminal paper of Zheng et al. (2018), which introduced the NOTEARS method, cleverly exploited the correspondence between the number of cycles in a matrix and the trace of that matrix’s power. A similar connection is now known to hold for other continuous functions and is exploited by methods such as DAGMA (Bello et al., 2022), which we use in this work, as well as those of Ng et al. (2020), Yu et al. (2021), and Gillot and Parviainen (2022). These approaches that use continuous characterizations of acyclicity are appealing since they scale to large graphs and can be augmented with discrete combinatorics (Manzour et al., 2021; Deng et al., 2023) if required. Though not pursued here, our projection onto DAG space is also amenable to discrete techniques.
Finally, although distributions over DAGs via projections is a new idea, others have used projections to construct distributions over non-graphical structures. Xu and Duan (2023) designed distributions over sparse regression coefficients by projecting a vector drawn from an arbitrary continuous distribution onto the $\ell_{1}$ ball. They illustrated several computational advantages of their distributions as priors in a Hamiltonian Monte Carlo (HMC) framework over classical spike-and-slab priors and prove that the resulting posterior enjoys the minimax concentration rate. Xu et al. (2023) extended this work to general proximal mappings with HMC, which includes the convex $\ell_{1}$ projection as a special case. While our approach projects matrices onto the $\ell_{1}$ ball, we also impose acyclicity, giving rise to a nonconvex combinatorial optimization problem that does not constitute a proximal mapping.
@@ 3.1 Description
To recap the introduction, we induce a distribution over weighted acyclic adjacency matrices $W$ as $W=\operatorname{pro}_{\lambda}(\tilde{W})$, where $\operatorname{pro}_{\lambda}$ is defined in (2), $\tilde{W}\sim\mathbb{P}$ is a continuous matrix, and $\lambda\sim\pi$ bounds the $\ell_{1}$ norm of $W$ [^2] The distribution $\pi$ can be a Dirac for a distribution over graphs of fixed sparsity (same number of edges) or an exponential for a distribution over graphs of varying sparsity. Since acyclic graphs have no self-loops (i.e., $W$ always has a zero diagonal), we can take $\mathbb{P}$ to be any continuous distribution over the off-diagonal elements of $\tilde{W}$ (e.g., a multivariate Gaussian), with the diagonal elements fixed as zero. The matrix $\tilde{W}$ can thus be interpreted as the weighted adjacency matrix of a simple directed graph, with the projection removing all cycles.
Besides removing cycles, the projection maps $\tilde{W}$ onto the $\ell_{1}$ ball of size $\lambda$. This part of the projection is necessary for sampling parsimonious graphs. Without this constraint, the projection would return the densest possible acyclic graph with $(p^{2}-p)/2$ edges. The advantage of constraining $W$ to the $\ell_{1}$ ball, compared with modeling $\tilde{W}$ as a Laplace distribution, is that the $\ell_{1}$ ball gives probability mass on exact zeros beyond that already given by the acyclicity constraint. It is known that Laplace distributions do not concentrate exactly on zero except at their mode (Park and Casella, 2008).
Conceptual interpretation
The resulting distribution $p(W)$ can be understood as a mix of continuous components and exact zeros. To elaborate, first observe that the nonzeros in $W$ may be indexed by a sparsity pattern $S\in\{0,1\}^{p\times p}$ (i.e., an unweighted acyclic adjacency matrix). The nonzeros are then $W_{S}$, where $W_{S}$ is the restriction of $W$ to $S$. Because the $\ell_{1}$ constraint shrinks the nonzeros towards zero, the continuous $W_{S}$ is a transformation $f_{\lambda}$ of $\tilde{W}_{S}$, i.e., $W_{S}=f_{\lambda}(\tilde{W}_{S})$, where $f_{\lambda}$ is implicit in the projection. This mixture interpretation suggests an interesting representation of $p(W)$:
$p(W)\equiv\sum_{i=1}^{m}p(W_{S_{i}}\mid S_{i})p(S_{i})=\sum_{i=1}^{m}p(f_{\lambda}(\tilde{W}_{S_{i}})\mid S_{i})p(S_{i}),$
where $S_{1},\dots,S_{m}$ is the finite collection of all acyclic adjacency matrices. We emphasize that this representation is purely conceptual and therefore we do not define or use the corresponding distributions over $S$, $W_{S}$ or $\tilde{W}_{S}$. However, it highlights that $p(W)$ forms a distribution over acyclic adjacency matrices $S$ without explicitly having to model the discrete, high-dimensional probability mass function $p(S)$. Moreover, as we later show, the projection can be reformulated as a continuous optimization problem, meaning one can sample from $p(W)$ without combinatorial complexity.
@@ 3.2 Properties
One might ask if $p(W)$ is a valid distribution. Theorem 1 addresses this question in the affirmative by showing that the projection is almost surely unique and measurable. The proof is in Appendix A.
Theorem 1.
@@
Theorem 1 states that, for any continuous $\tilde{W}\sim\mathbb{P}$, the function $\operatorname{pro}_{\lambda}(\tilde{W})$ is uniquely defined and measurable. This important result implies that a projected distribution is a valid distribution since a measurable function $g$ applied to a random variable $z$ produces a well-defined random variable $g(z)$.
@@ 4.1 Continuous reformulation
To facilitate scalability, we reformulate the combinatorial projection (2) as an equivalent continuous optimization problem. This recharacterization follows recent work (Zheng et al., 2018; Bello et al., 2022) that established a useful and important equivalence relationship for certain functions $h(W)$:
$h(W)=0\iff W\in\mathrm{DAG}.$
Bello et al. (2022) showed that for all matrices $W\in\mathbb{W}:=\{W\in\mathbb{R}^{p\times p}:\rho(W\circ W)<1\}$, where $\rho$ is the spectral radius and $\circ$ is the Hadamard product, the function $h(W):=-\log\,\det(I-W\circ W)$ satisfies this property. This function, which constitutes the main innovation of DAGMA, attains the value zero if and only if there are no cycles in $W$. Thus, its level set at zero exactly corresponds to the entire set of DAGs. Though it is a nonconvex function, $h(W)$ is useful because it is both continuous and differentiable in $W$. We thus exploit this equivalence of constraints and rewrite projection (2) as
@@
As we discuss next, it is possible to (i) solve this projection using first-order methods and (ii) evaluate its gradients analytically. These attributes pave the way for scalable, gradient-based posterior learning.
@@ 4.2 Algorithms
To solve projection (3), we employ a two-step procedure that produces a solution on the intersection of sets. Step 1 projects the matrix $\tilde{W}$ onto the set of acyclic matrices. We call the result $\hat{W}$. Step 2 projects $\hat{W}$ onto the $\ell_{1}$ ball, yielding the final output $W$ that remains acyclic but is further sparse.
For the acyclicity projection in Step 1, we use a path-following algorithm similar to Bello et al. (2022). This algorithm reformulates the problem by shifting the acyclicity constraint into the objective: